if i need to mix 5 mls per every 4 litres
how many mls would i mix for a 5 gallon bucket

Answer by stanbon(75887)

if i need to mix 5 mls per every 4 litres
how many mls would i mix for a 5 gallon bucket
-----
Each gallon is 3.785 liters
---
5 gallons is 5*3.785 = 18.925 liters
----
# mls needed = 5*(18.925/4) = 23.66 ml
==================
Cheers,
Stan H.
==================